OpenStudy (anonymous):

what percentage is 15 of 33, exactly?

OpenStudy (anonymous):

15/33 = 5/11 = 0.45454545... So there is no finite decimal representation

OpenStudy (uh60blackhawk):

just divide 15/33 and it will equal to 0.45